# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ Abstract.
This paper investigates the neural dropout method as a post-processing bias mitigation for deep neural networks (DNNs). Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications. While neural networks are exceptionally good at finding statistical patterns from data, they are notorious for over-fitting the training datasets that may encode and amplify existing biases from the historical data. Existing bias mitigation algorithms often require either modifying the input dataset or modifying the learning algorithms. We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve fairness of pre-trained DNNs. However, finding the ideal set of neurons to drop is a combinatorial problem.
We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference after training. Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility. We show that our design of randomized algorithms are effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation. We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results. Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
@@ 1. Introduction
Despite many advances made possible by AI, there are challenges that require understanding the dimensions and implications of deploying AI-driven software solutions. One such concern about the trustworthiness of AI is discrimination. Unfortunately, there are plenty of fairness defects in real systems. Parole decision-making software was found to harm black and Hispanic defendants by falsely predicting a higher risk of recidivism than for non-Hispanic white defendants (Julia Angwin and Kirchne, 2021); Amazon’s hiring algorithm disproportionately rejected more female applicants than male applicants (Ingold and Soper, 2016); and data-driven auditing algorithm selected black taxpayers with earned income tax credit claims (EITC) at much higher rates than other racial groups for an audit (Tankersley, 2023). As evidenced by these examples, resulting software may particularly disadvantage minorities and protected groups and be found non-compliant with law such as the US Civil Rights Act (Blumrosen, 1978). Hence, helping programmers and users to mitigate unfairness in social-critical data-driven software systems is crucial to ensure inclusion in our modern, increasingly digital society.
@@
Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model). However, when the decision logic of AI is encoded via deep neural networks, it becomes increasingly challenging to mitigate unfairness due to its black-box uninterpretable nature.
Since the DNN-based software development is derived by principle of information bottleneck (Tishby et al., 2000), which provides a trade-off between over-fitting and parsimonious signals; we posit that over-fitted DNN models may contribute and amplify historical biases from the training datasets. Therefore, prevalent techniques such as dropout methods (Srivastava et al., 2014; Srivastava, 2013)—a process of randomly dropping neurons during training—might be effective in mitigating unfairness in the pre-trained DNN models. While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods for mitigating unfairness in pre-trained DNN models, as a post-processing mitigation method.
@@
We evaluate NeuFair over $7$ deep neural network benchmarks as trained over $5$ socially critical applications with significant fairness implications. We found that randomized strategies of neuron drops can successfully improve fairness (up to 69%) with a minimal utility degradation in the vast majority of cases. We also report a pathological case and reasons behind a failure of NeuFair. We also observe that NeuFair can improve fairness and utility together and provide intuitive explanations of such phenomena. Furthermore, we examine different hyperparameter configuration options of randomized algorithms. While some hyperparameters always influence fairness with positive or negative impacts; we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark. Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023). In summary, the key contributions of this paper are:
@@ 3.2. Inference Time Dropout for Fairness
Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of neural networks by preventing overfitting. The key idea is that randomly dropping out neurons in a layer during training can help prevent overfitting of the neural network. Dropout randomly sets all $w_{i}$ to 0 for each neuron in the hidden layers with probability $p$ during training. Once training is complete, dropout is not used, and all the neurons in the network are utilized to make predictions. While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the model during inference after training can significantly improve fairness with a minimal impact on model performance. However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
@@
A brute-force search would be exponential in the size of DNN as we have $2^{|N|}$ possible subsets. The running time becomes prohibitively expensive, even for small neural networks. We define desirable states as those states that have a good fairness-performance tradeoff. To find the desirable subset of neurons, we explore different types of randomized algorithms to improve fairness via model inference dropout.
@@ 4.1. Simulated Annealing Search
We formulate the problem of finding the desirable subset of neurons as a discrete optimization problem and solve it using Simulated Annealing. Simulated Annealing (SA) (Bertsimas and Tsitsiklis, 1993) is a probabilistic algorithm that finds the global minima (w.r.t some cost function) in a large search space with high probability. Algorithm 1 presents a generic template to apply SA to a search space optimization problem. Figure 1 overviews the steps in our bias mitigation approach with the simulated annealing search. We now define the core concepts used in our SA algorithm.
@@ 5.1. Datasets and Models
We used DNNs as the machine learning model, trained over the dataset benchmarks. Table 1 highlights the architectures of the DNN models used for each dataset. We use ReLU as the activation function after the linear layers. For the Bank, Default, and Compas datasets we use a dropout of 0.2 during training. For the Adult and MEPS16 datasets, we set the dropout to 0.1. For the Compas dataset, we use Adam as the optimizer with a learning rate of $0.001$. For the other datasets, we use SGD with a learning rate of $0.01$. For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features. For the Compas dataset, we use a version of the dataset that has 12 features after feature selection that was used in (Tizpaz-Niari et al., 2022). The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
@@ 5.3. Experimental Setup
For each dataset, we evaluate the performance of the NeuFair algorithms using $10$ different random seeds. The seeds are set for torch, numpy, and scikit-learn libraries before DNN training. Specifically, the seed determines the randomness of the training, validation, and test splits and the randomness of model training, such as sampling batches and initializing the weights of the DNN. The NeuFair algorithms themselves are not seeded. Each SA and RW run is unique.
During training, the training dataset was used for gradient descent, and the validation dataset was used to monitor the F1 score and determine the hyperparameters. We use the model from the epoch with the highest validation dataset. The same validation dataset is used by NeuFair to identify the desirable set of neurons to drop. Finally, we evaluate the fairness of the original and fixed (repaired) models using the test dataset.
@@
The F1 threshold is 98% of the validation F1 score of the unfair neural network, i.e., we tolerate a 2% degradation in the F1 score to improve fairness. A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold. For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons depending on the overall size of the network. The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively. For all datasets, we use a train/validation/test split of 60%/20%/20%.
Table 2. Datasets used in our experiments
@@
Table 3. Fairness and model utility metrics of the unrepaired (original) neural network.
@@
Table 4. Fairness and model utility metrics of the repaired neural network.
@@ 5.3.1. Effectiveness of randomized algorithms on mitigating unfairness using dropout (RQ1)
Figure 2 and Figure 3 highlight the cost of the best state found for the SA and RW over time, respectively, for 10 different seeds per data subject using the logarithmic scale. In all cases, we see a positive trend where the cost improves over time, and the search identifies better subsets of neurons to drop. We note the seemingly flat plots correspond to those cases where a good state was found in the beginning. The exception is the black line in the sub-plot for the Default dataset in both figures as it corresponds to a failed run; i.e., the best state found has an F1 score lesser than 0.98×0.98$\times$0.98 × the F1 score of the initial state. Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases. In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$. The progress graphs of the RW strategy are “steep” initially and rapidly improve as the initial states are unfair and also heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold. However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability. So, the progress graphs are not as steep initially.
Table 3 and Table 4 highlight the mean and $95\%$ confidence intervals for the F1 score, accuracy, and EOD for all data subjects aggregated across 10 seeds. We observe that in all cases, both NeuFair algorithms improve the EOD. However, the SA algorithm outperformed RW in terms of EOD in all cases for the validation set while ensuring that the validation F1 score reduces by at most $2\%$. We can conclude that SA is the more effective algorithm in identifying the desirable state during the search procedure, subject to the F1 threshold. On the test set, we observe that both NeuFair algorithms improve the EOD. SA outperforms RW for 4 datasets (Adult (Sex), Bank, Default, and MEPS16). We observe the largest absolute improvement of $12.215\%$ in EOD ($20.641\%$ to $8.426\%$ ) for the MEPS16 dataset using the SA algorithm. The largest relative improvement of $69.43\%$ ($8.962\%$ to $2.749\%$) was observed in the Default dataset using SA. However, SA did not improve the EOD on the COMPAS (Sex), which is a pathological case.
@@ 5.3.2. Dropout strategies improving both fairness and utility (RQ2)
The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is negative class and 1 is positive) as more and more neurons are dropped out from the DNN. The increase in negative predictions favors the true negatives as the datasets are higly imbalanced with more negative samples. The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class. The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class. The accuracy increases as the increase in true negatives outweighs the drop in true positives. The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and false negatives increase.
@@ 5.3.3. Hyperparameters of randomized algorithms and fairness (RQ3)
  • The F1 penalty multiplier has a more nuanced effect on the unfairness as it controls the exploration vs. exploitation trade-offs. A low penalty multiplier increases the probability of state transitions when the F1 score is less than our threshold, increasing the SA run’s randomness to explore more of the search space outside of the current best state. A high penalty multiplier would decrease the probability of state transitions, thereby encouraging the SA algorithm to exploit the search space near the current best state. The penalty multiplier ‘p’ directly affects the cost difference $\Delta E$ between states with acceptable and unacceptable F1 scores. From Line 14 in Algorithm LABEL:alg:neuron_repair, we can see that a higher cost difference results in a lower transition probability. For a fixed temperature, a higher ‘p’ would increase the cost difference between good and bad states, thereby decreasing the probability of transitioning and encouraging exploitation. A lower ‘p’ would result in a lower cost difference and increase the transition probability, thereby encouraging exploration. To determine the effect of the penalty multiplier, we run experiments with the Adult, MEPS16, Default, and Bank datasets with one seed by varying the F1 penalty multiplier $p\in\{0.5,1.0,1.5,2.0,2.5,3.0\}$ and F1 threshold multiplier to 0.98. Figure 4 shows the results. The red dot in the Default dataset corresponds to a run in which SA failed to find a desirable state within the F1 threshold. We can see that Adult dataset performs better on lower F1 penalties. The Bank dataset has a “sweet” spot in between while Default performs worse on intermediate F1 penalties. The MEPS16 does not follow a clear trend. The results show that the F1 penalty multiplier is a hyperparameter that trades off explorations vs. exploitation and requires tuning for each dataset and model.
@@ 6. Discussion
Effectiveness of Randomized Algorithms. To understand why randomized algorithms are effective in mitigating unfairness, we compare them to brute-force strategies. We create a small DNN for the Adult (Sex) dataset with 2 hidden layers, each with 16 neurons, to allow the brute force search to explore the entire space. We use a restricted search space with $n_{l}=4$ and $n_{u}=9$ that has a total of $43,076,484$ states. The time-out for SA runs is one hour. The brute force runs take 60 hours to explore the entire search space. Table 7 highlights the results of both search strategies. SA finds the global optimal solution for two seeds in the search space. The EOD Delta column highlights the difference in cost of the states identified by SA and brute force. The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states. SA is faster in reaching a good state compared to brute force as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost. Another reason for SA’s effectiveness is that it uses a probabilistic transition function that can transition to a state that has a worse cost and escape local minima. Gradient-based methods often struggle with the problem of getting stuck in local minima. Table 8 shows the distribution of the Best, Good, and Bad states. The best states are the states with the global optimal cost. The cost difference between the good and best states is less than 0.05. Bad states are states whose F1 scores are less than the 98% of the baseline F1 score. We observe that the best and good states are a negligible amount of the search space. On the other hand, over 90% of the states are bad states. This suggests that the good states occupy small “pockets” of the search space. The cost gradient leads SA to effectively identify such pockets. The probabilistic transition function enables SA to explore the region without getting stuck in local minima.
@@
Threat to Validity. To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines and repeat the experiments 10 times, reporting both the average and 95% confidence intervals. In addition, we study the progress of algorithms during their runs, not just the final outcomes. To ensure that our results are generalizable and address external validity, we perform our experiments on seven DNN benchmarks with various architectures. While similar architectures have been used in the prior works (Monjezi et al., 2023; Zhang et al., 2020b; Zheng et al., 2022; Zhang et al., 2021), it is an open problem whether these datasets and DNN models are sufficiently representative for showing the effectiveness of NeuFair.

### v2

@@ Abstract.
This paper investigates neuron dropout as a post-processing bias mitigation for deep neural networks (DNNs). Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications. While neural networks are exceptionally good at finding statistical patterns from data, they may encode and amplify existing biases from the historical data. Existing bias mitigation algorithms often require modifying the input dataset or the learning algorithms. We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve the fairness of pre-trained DNNs. However, finding the ideal set of neurons to drop is a combinatorial problem.
We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference after training. Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility. We show that our design of randomized algorithms is effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation. We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results. Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
@@ 1. Introduction
Despite many advances made possible by AI, some challenges require understanding the dimensions and implications of deploying AI-driven software solutions. One such concern about the trustworthiness of AI is discrimination. Unfortunately, there are plenty of fairness defects in real systems. Parole decision-making software was found to harm black and Hispanic defendants by falsely predicting a higher risk of recidivism than for non-Hispanic white defendants (Julia Angwin and Kirchne, 2021); Amazon’s hiring algorithm disproportionately rejected more female applicants than male applicants (Ingold and Soper, 2016); and data-driven auditing algorithm selected black taxpayers with earned income tax credit claims (EITC) at much higher rates than other racial groups for an audit (Tankersley, 2023). As evidenced by these examples, resulting software may particularly disadvantage minorities and protected groups and be found non-compliant with laws such as the US Civil Rights Act (Blumrosen, 1978). Hence, helping programmers and users to mitigate unfairness in social-critical data-driven software systems is crucial to ensure inclusion in our modern, increasingly digital society.
@@
Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model). However, when the decision logic of AI is encoded via deep neural networks, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
We posit that a subset of neurons in a neural network disparately contributes to unfairness. Removing these neurons during inference as a post-processing operation on a trained DNN can improve fairness. Therefore, prevalent techniques such as dropout methods (Srivastava et al., 2014; Srivastava, 2013)—a process of randomly dropping neurons during training—might be effective in mitigating unfairness in pre-trained DNN models. While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods as a post-processing bias mitigation method for improving the fairness of pre-trained DNN models.
@@
We evaluate NeuFair over $7$ deep neural network benchmarks trained over $5$ socially critical applications with significant fairness implications. We found that NeuFair can improve fairness (up to 69%) with minimal utility degradation in most cases. We also report a pathological case and reasons behind a failure of NeuFair. We also observe that NeuFair can simultaneously improve fairness and utility and provide intuitive explanations of such phenomena. Furthermore, we examine different hyperparameter configuration options of randomized algorithms. While some hyperparameters always influence fairness with positive or negative impacts; we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark. Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023).
In summary, the key contributions of this paper are:
@@ 3.2. Inference Time Dropout for Fairness
Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of DNNs by preventing overfitting. Dropout sets all $w_{i}$ to 0 for a random set of neurons in the hidden layers with some probability during training. Once training is complete, dropout is not used, and all the neurons in the DNN are utilized to make predictions. While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on model performance. However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
@@
We define desirable states as those states that have a good fairness-performance tradeoff. A brute-force search to find desirable neurons is exponential in the size of DNN as we have $2^{|N|}$ possible subsets. The running time becomes prohibitively expensive, even for small DNNs. To find the desirable subset of neurons, we explore different types of randomized algorithms to improve fairness via model inference dropout.
@@ 4.1. Simulated Annealing Search
We formulate the problem of finding the desirable subset of neurons as a discrete optimization problem and solve it using Simulated Annealing. Simulated Annealing (SA) (Bertsimas and Tsitsiklis, 1993) is a probabilistic algorithm that finds the global minima (w.r.t some cost function) in a large search space with high probability. Algorithm 1 presents a generic template to apply SA to a search space optimization problem. Figure 1 overviews the steps in our bias mitigation approach with the SA search. We now define the core concepts used in our SA algorithm.
@@ 5.1. Datasets and Models
We used DNNs trained over the dataset benchmarks as the machine learning model. Table 1 highlights the architectures of the DNN models used for each dataset. We use ReLU as the activation function after the linear layers. For the Bank, Default, and Compas datasets we use a dropout of 0.2 during training. For the Adult and MEPS16 datasets, we set the dropout to 0.1. For the Compas dataset, we use Adam as the optimizer with a learning rate of $0.001$. For the other datasets, we use SGD with a learning rate of $0.01$. For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features. For the Compas dataset, we use a version of the dataset that has 12 features after feature selection that was used in (Tizpaz-Niari et al., 2022). The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
@@ 5.3. Experimental Setup
For each dataset, we evaluate the performance of the NeuFair algorithms using $10$ different random seeds. The seeds are set for the torch, numpy, and scikit-learn libraries before DNN training. Specifically, the seed determines the randomness of the training, validation, and test splits and the randomness of model training, such as sampling batches and initializing the weights of the DNN. The NeuFair algorithms themselves are not seeded. Each SA and RW run is unique.
During training, the training dataset was used for gradient descent, and the validation dataset was used for hyperparameter tuning. We use the model from the epoch with the highest validation F1 score. The validation dataset is also used by NeuFair to identify the desirable set of neurons to drop. We evaluate the fairness of the original and fixed (repaired) models using the test dataset.
@@
The F1 threshold is 98% of the validation F1 score of the unfair DNN, i.e., we tolerate a 2% degradation in the F1 score to improve fairness. A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold. For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons depending on the overall size of the network. The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively. For all datasets, we use a train/validation/test split of 60%/20%/20%. We set the threshold for the acceptance probabilities in the temperature initialization algorithm (Ben-Ameur, 2004) to $0.75$.
Table 2. Datasets used in our experiments.
@@
Table 3. Fairness and model utility metrics of the unrepaired (original) DNN.
@@
Table 4. Fairness and model utility metrics of the repaired DNN.
@@ 5.3.1. Effectiveness of randomized algorithms on mitigating unfairness using dropout (RQ1)
Figure 2 and Figure 3 highlight the cost of the best state found for the SA and RW over time, respectively, for 10 different seeds per data subject using the logarithmic scale. In all cases, we see a positive trend where the cost improves over time, and the search identifies better subsets of neurons to drop. We note the seemingly flat plots correspond to those cases where a good state was found in the beginning. The exception is the black line in the sub-plot for the Default dataset in both figures as it corresponds to a failed run; i.e., the best state found has an F1 score lesser than 0.98×0.98$\times$0.98 × the F1 score of the initial state. Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases. In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$. The progress graphs of the RW strategy are “steep” initially and as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold. However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability. So, the progress graphs are not as steep initially.
Table 3 and Table 4 highlight the mean and $95\%$ confidence intervals for the F1 score, accuracy, and EOD for all data subjects aggregated across 10 seeds. We observe that in all cases, both NeuFair algorithms improve the EOD. However, SA outperformed RW in all cases for the validation set while ensuring that the F1 score reduces by at most $2\%$. We can conclude that SA is the more effective algorithm in identifying the desirable state during the search procedure, subject to the F1 threshold. On the test set, we observe that both NeuFair algorithms improve the EOD. SA outperforms RW for 4 datasets (Adult (Sex), Bank, Default, and MEPS16). We observe the largest absolute improvement of $12.215\%$ in EOD ($20.641\%$ to $8.426\%$ ) for the MEPS16 dataset using SA. The largest relative improvement of $69.43\%$ ($8.962\%$ to $2.749\%$) was observed in the Default dataset using SA. However, SA did not improve the EOD on the COMPAS (Sex), which is a pathological case.
@@ 5.3.2. Dropout strategies improving both fairness and utility (RQ2)
The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is the negative class and 1 is positive) as neurons continue to be dropped out from the DNN. The increase in negative predictions favors the true negatives as the datasets are highly imbalanced with more negative samples. The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class. The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class. The accuracy increases as the increase in true negatives outweighs the drop in true positives. The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and false negatives increase.
@@ 5.3.3. Hyperparameters of randomized algorithms and fairness (RQ3)
  • The F1 penalty multiplier has a more nuanced effect on the unfairness as it controls the exploration vs. exploitation trade-offs. A low penalty multiplier increases the probability of state transitions when the F1 score is less than our threshold, increasing the SA run’s randomness to explore more of the search space outside of the current best state. A high penalty multiplier would decrease the probability of state transitions, thereby encouraging the SA algorithm to exploit the search space near the current best state. The penalty multiplier p directly affects the cost difference $\Delta E$ between states with acceptable and unacceptable F1 scores. From Line 14 in Algorithm LABEL:alg:neuron_repair, we can see that a higher cost difference results in a lower transition probability. For a fixed temperature, a higher p would increase the cost difference between good and bad states, thereby decreasing the probability of transitioning and encouraging exploitation. A lower p would result in a lower cost difference and increase the transition probability, thereby encouraging exploration. To determine the effect of the penalty multiplier, we run experiments with the Adult, MEPS16, Default, and Bank datasets with one seed by varying the F1 penalty multiplier $p\in\{0.5,1.0,1.5,2.0,2.5,3.0\}$ and F1 threshold multiplier to 0.98. Figure 4 shows the results. The red dot in the Default dataset corresponds to a run in which SA failed to find a desirable state within the F1 threshold. We can see that Adult dataset performs better on lower F1 penalties. The Bank dataset has a “sweet” spot in between while Default performs worse on intermediate F1 penalties. The MEPS16 does not follow a clear trend. The results show that the F1 penalty multiplier is a hyperparameter that trades off explorations vs. exploitation and requires tuning for each dataset and model.
@@ 6. Discussion
Effectiveness of Randomized Algorithms. To understand why randomized algorithms are effective in mitigating unfairness, we compare them to brute-force strategies. We create a small DNN for the Adult (Sex) dataset with 2 hidden layers, each with 16 neurons, to allow the brute force search to explore the entire space. We use a restricted search space with $n_{l}=4$ and $n_{u}=9$ that has a total of $43,076,484$ states. The time-out for SA runs is one hour. The brute force exhaustive search takes 60 hours. Table 7 highlights the results of both search strategies. SA finds the global optimal solution for two seeds in the search space. The EOD Delta column highlights the difference in cost of the states identified by SA and brute force. The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states. SA is faster than brute force in reaching a good state as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost. Additionally, SA uses a probabilistic transition function that can transition to worse states and potentially escape local minima. Gradient-based methods often get stuck in local minima. Table 8 shows the distribution of the Best, Good, and Bad states. The best states are the states with the global optimal cost. The cost of the good states is within 0.05 of the optimal cost. Bad states are states whose F1 scores are less than the 98% of the baseline F1 score. We observe that the best and good states are a negligible amount of the search space. On the other hand, over 90% of the states are bad states. This suggests that the good states occupy small “pockets” of the search space. The cost gradient effectively leads SA to identify such pockets and the probabilistic transitions prevent it from getting stuck in local minima.
@@
Threat to Validity. To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines and repeat the experiments 10 times, reporting both the average and 95% confidence intervals. In addition, we study the progress of algorithms during their runs, not just the outcomes. To ensure that our results are generalizable and address external validity, we perform our experiments on seven DNN benchmarks with various architectures. While similar architectures have been used in the prior works (Monjezi et al., 2023; Zhang et al., 2020b; Zheng et al., 2022; Zhang et al., 2021), it is an open problem whether these datasets and DNN models are sufficiently representative for showing the effectiveness of NeuFair.